# Joint Liability Model with Adaptation to Climate Change

### v1

@@ 3.1 Model Assumptions
Prior to constructing the model, it is necessary to state the following assumptions:
  1. 1.
@@
    The borrowers (who are the individual farmers) produce an agricultural product. For each borrower, the probability of high production (i.e., a successful project) is the same as $e$ and the probability of low production (i.e., a failed project) is $1-e$ respectively.
  2. 2.
@@
    A higher individual-ESE score means the farmer has a stronger ability to succeed in his/her project due to the high-quality comprehensive evaluation. The probability of high yield is assumed to be linearly related to the borrower’s $ESE$ score; that is we assume that there is a linear relationship between $e$ and $ESE$ score: $e=k\cdot ESE+b$. The parameter $k$ represents the weight or strength of the impact of the borrower’s ESE score on the probability of high yield. A higher value of $k$ indicates that a higher ESE score leads to a more significant positive impact on the probability of success. This aligns with the idea that a borrower’s positive environmental, social, and economic practices contribute to project success. In other words, higher $ESE$ scores suggest better alignment with responsible practices. The constant $b$ represents the baseline or constant contribution of factors beyond an individual’s control, such as climatic and geographical conditions. Even if a borrower has a lower $ESE$ score, this constant term acknowledges that external factors can still influence the project’s success. It’s essential to account for these factors to avoid unfairly penalizing borrowers for circumstances beyond their control.
@@
    The selection of the linear equation $e=k\cdot ESE+b$ is rooted in both the characteristics of the variables involved and a keen understanding of practical applicability. Notably, the non-negativity of $k$, $b$, and $ESE$ ensures that the resulting probability, denoted as $e$, maintains its intrinsic non-negative nature. As $ESE$ encompasses a continuous range from 0 to 100, it inherently aligns with this non-negative property. It is important to stress that the intuitive requirement for $e$ to lie between 0 and 1, which is a fundamental property of a probability, is elegantly met by setting $b=0$ and $k=\frac{1}{100}$. By adhering to these values, the probability of $e$ adheres to the established probability range. In addition, the model accommodates real-world scenarios by allowing for a non-zero baseline influence ($b>0$), often linked to exogenous factors. To this end, a modest value such as $b=0.3$ can be accommodated without violating the probability constraints. Notably, this adjustment merely necessitates setting $k=\frac{0.7}{100}$, thereby guaranteeing that $e$ remains within the bounds of a probability. While it’s acknowledged that $e$ cannot span the interval from 0 to 0.3, this limitation is justified by underscoring $e$ as a representation of the probability of high yield, which is inherently a positive entity in practical settings. Hence, the rationale underlying the proposed equation effectively marries mathematical coherence with the nuanced demands of reality.
    The exploration of using a sigmoid function to model $e$ as $e=\frac{1}{1+\exp^{-ESE}}$ or $e=\frac{1}{1+\exp^{-(k\cdot ESE+b)}}$ is an intriguing avenue to consider. However, upon a closer examination, a critical observation emerged: under the assumption that $k$, $b$, and $ESE$ are non-negative, the sigmoid-based approach resulted in a constraint where $e$ precludes the values below $1/2$. This deviation from the alignment with realistic scenarios and the entire spectrum of probability values was deemed undesirable. Thus, while the sigmoid formulation may possess merits in other contexts, its inherent restriction on precluding the probability values below $1/2$ is at odd with the overarching goal of accurately representing the probability of high yield across its entire spectrum. As a result, the decision was made to retain the linear equation $e=k\cdot ESE+b$, which adheres to both the established criteria and delivers the sought-after flexibility in representing diverse probability values.
  3. 3.
@@
    The amount borrowed by each borrower is denoted by $L$, and the amount for this borrower to be repaid is denoted by $w$. The unit selling price of the product is denoted by $p$, and the total income of high production is denoted by $p\overline{Y}$, while the total income of low production is denoted by $p\underline{Y}$.
  4. 4.
@@
    High-yield borrowers repay their loans and low-yield borrowers pay back all the income they earn. As long as this group repays the total loans, this group can continue to get the subsequent loans. In the event that all borrowers are unsuccessful in their projects, the group will not be eligible for any future loans. This may result in a negative credit history for all members of the group, which could lead to further restrictions such as being unable to use public transportation or purchase high-cost items. Therefore, if all borrowers fail, they may have to use all of their earnings to avoid receiving negative credit reports, even though their total earnings will be less than the total repayments.
  5. 5.
@@
    To increase the successful probability $e$, the borrower incurs additional costs such as purchasing green pesticides, advanced tools, etc. A convex disutility function, denoted by $C(e)$, is included in the specification of the total utility function.
@@ 3.2 Joint Liability Contracts with a Group Size of Two
Under the joint liability agreement, each farmer’s ex-ante expected profit $\pi_{i}$ can be written as
@@
To clarify, both farmers have a chance of achieving a successful harvest $\overline{Y}$ with a probability of $e^{2}$, which gives them the profit of $p\overline{Y}-w$. The probability of one farmer defaulting is $e(1-e)$, which means that farmer $i$ would receive her own surplus from a successful harvest minus the other farmer’s deficit, $p\overline{Y}+p\underline{Y}-2w$. If both producers fail at the same time, they will pay all of their incomes to the lenders, which means the profit is zero. The detailed profit distributions can be summarized in Table 3.
@@
The farmer succeeds, another farmer succeeds too. $e^{2}$ $p\overline{Y}-w$
The farmer succeeds, another farmer fails. $e(1-e)$ $p\overline{Y}+p\underline{Y}-2w$
The farmer fails, another farmer succeeds. $(1-e)e$ $p\underline{Y}-p\underline{Y}=0$
The farmer fails, another farmer fails. $(1-e)^{2}$ $p\underline{Y}-p\underline{Y}=0$
@@
Before setting the total objective (utility) function for the whole group, we first specify the following constraints for the optimization program.
  1. 1.
    If the group wants a subsequent loan from the lender, the group’s total repayment amount must be affordable even in the worst-case scenario when everyone else’s project has failed except his/hers:
@@
  2. 2.
    Strategic default in a joint liability contract occurs when one or more parties intentionally chooses or choose to default on their obligations, even though they have the ability to pay. This is often done to gain a strategic advantage, such as forcing the other parties to pay a greater share of the debt or renegotiating the terms of the contract. To avoid a strategic default for each successful borrower, the payoff of strategically defaulting cannot be larger than the payoff of repaying and being refinanced:
    1. (a)
      Assuming the other borrower also succeeds, this borrower will only be responsible for their own payment. In such a scenario, the earnings plus the discounted cost of the next loan would at least be equal to the earnings of strategically defaulting.
@@
    2. (b)
      If the other borrower fails, then this borrower needs to pay for himself/herself as well as part of the other borrower’s repayment amount. Now, the earnings plus the discounted cost of the next loan would also at least equal the earnings of strategically defaulting.
@@
  3. 3.
    The lender must be able to sustain the lending game over the periods or at least break even. Thus, the expected repayment amount of each borrower has to be at least as large as $L(1+\epsilon)$, where $\epsilon$ is a risk-free rate. Given that all constraints have been established under the presumption that the group will successfully repay the total loans, to ensure their eligibility for subsequent loan rounds, we have not factored in the scenario where all borrowers within the group experience failure on the LHS of Constraint 4.
@@
Since $p\underline{Y}-w<0$, it is easy to see that constraint 3 is tighter than constraint 2. Hence, we will only consider constraint 1, constraint 3, and constraint 4. Firstly, $\pi_{i}$ and $w$ have a strictly negative relationship as can be observed from Equation 3.1, then if we want to maximize $\pi_{i}$, $w$ has to be set at a minimum value. From constraint 4, the minimum value of $w$ should be set at
@@
Next, we can substitute $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into constraints 1, and then it can be rewritten as:
@@
and the first loan ceiling bound can be expressed in Equation 3.8 as:
@@
After substituting $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into constraints 3, it can be rewritten as:
@@
and thus we can obtain another bound of the loan ceiling as in Equation 3.11:
@@
Now, we substitute $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into the objective function 3.1 and replace $e=k\cdot E+b$. The ex-ante expected profit function of an individual borrower (farmer) can be rearranged as
@@
To simplify the problem further, we set $C(e)=\frac{1}{2}ce^{2}$ (where the cost parameter $c$ is used to scale the total cost) as the disutility function. The disutility function $C(e)=\frac{1}{2}ce^{2}$ is a commonly used function in economics to represent the cost or disutility associated with effort or work. See e.g. Clark and Oswald (1996) and Kahneman et al. (1997). The function is quadratic, meaning that as effort increases, the cost or disutility of that effort increases at an increasing rate. In the context of an individual agricultural farmer, it is reasonable to use a disutility function that acknowledges that while some effort may be necessary and satisfying, there is a point at which additional effort becomes more and more unpleasant or burdensome. For instance, a farmer may be willing to put in extra effort to cultivate crops or tend to livestock, but as the required effort increases, the farmer may experience fatigue, stress, and reduced productivity, which can negatively impact their overall well-being. Furthermore, the quadratic form of the disutility function is also supported by empirical evidence on the relationship between effort and well-being. It is reasonable to assume that people tend to experience a phenomenon known as a diminishing marginal utility of effort, which implies that the additional unit of well-being gained by exerting more effort decreases as the level of effort increases.
3.3 Some Propositions
Now that we have listed the model assumptions and set up for our analysis, we can move on to exploring key propositions of this model. These propositions provide insights into the optimal individual-ESE scores, optimal group size, and loan ceiling constraints.
Proposition 1: Find the optimal individual-ESE score.
Finding the optimal ESE score for both lenders and borrowers in a joint liability problem is critical because it can help address the challenge of promoting sustainability in agricultural production while ensuring access to credit for small-scale farmers. By incorporating environmental and social factors into credit evaluation systems, it can incentivize farmers to adopt sustainable practices that can lead to positive environmental outcomes and long-term economic benefits. At the same time, lenders can mitigate the risks associated with lending to farmers by incorporating sustainability factors into their credit evaluation process.
To find the optimal individual-ESE score (noted in formulas as $E$) score, we take the first-order partial derivative w.r.t $E$ of the ex-ante expected payoff function to obtain the following result:
@@ 3.3 Some Propositions
After rearranging of terms, we can get the optimal individual-ESE score ($E$ is between 0 and 100) as
@@
Proposition 2: Reduce the two constraints about loan ceiling to only one constraint.
The two loan ceiling constraints (constraint 1 and constraint 3) arise from analyzing the success or failure of the other co-loaners assuming our own success, but most of the previous studies in this area have not explored the connection between these constraints. For instance, in Rezaei et al. (2017), they found that a feasible range of loan ceiling depends on the borrower’s discount rate $\delta$. Nevertheless, it is possible to reduce these two constraints to a single constraint in our model.
Recall the two constraints about the loaning ceiling in Equation 3.8 and Equation 3.11, named respectively as $L_{1}$ and $L_{2}$.
@@
Next, we show that $L_{1}$ is always larger than $L_{2}$, which means that we can only use $L_{2}$ as the final loan ceiling bound. It simplifies the problem and potentially makes it easier to analyze and solve the model. In addition, $L_{2}$ takes $\delta$ into consideration, which may include more information and additional key factors. This conclusion can also be illustrated by Figure 1.
@@
Proof:
If $L_{1}\leqslant L_{2}$, then
@@
After rearranging the terms, we get:
@@
From the above results, if $L_{1}\leqslant L_{2}$, the discount factor $\delta$ is always greater than 1, which leads to a contradiction. So, we conclude that $L_{1}>L_{2}$.
Proposition 3: A higher individual-ESE score leads to a higher loan ceiling.
Next, we examine the relationship between the ESE score and the loan ceiling. Intuitively, a borrower with a high-quality ESE score should be eligible for a higher loan amount, and therefore the loan ceiling should also be higher.
Since $L_{1}$ is always greater than $L_{2}$, we will use $L_{2}$ to determine the loan ceiling, which defines the maximum value of $L_{2}$ to be the loan ceiling.
@@
Then, we take a partial derivative of $L_{2}$ w.r.t. $E$ to get:
@@
Obviously, the denominator is greater than 0. The numerator is also greater than 0, which can be illustrated by writing out the following result:
@@
A borrower with a higher optimal ESE score, which means he/she has better overall performance, will be perceived as less risky by the lender. As a result, the borrower will be offered a larger loan ceiling, which is the maximum amount of loan that a lender is willing to provide to a borrower. In turn, if a borrower wants to obtain a larger loan amount, he/she may be motivated to improve the ESE performance.

### v2

@@ 2.1 Model Assumptions
Prior to constructing the individual ESE–joint liability model, we state a number of assumptions underlying it.
@@
ASSUMPTION 1. The borrowers (who are the individual farmers) produce an agricultural product. For each borrower, the probability of high production (i.e., a successful project) is the same as given by $e$ and the probability of low production (i.e., a failed project) is given by $1-e$ respectively.
@@
ASSUMPTION 2. A higher individual–ESE score means that the farmer has a stronger ability to succeed in their project due to a high-quality comprehensive evaluation. For simplicity the probability of high yield is assumed to be linearly related to the borrower’s $ESE$ score; i.e., we assume that there is a linear relationship between $e$ and $ESE$ score: $e=k\cdot ESE+b$. The parameter $k$ represents the weight or strength of the impact of the borrower’s $ESE$ score on the probability of high yield. A higher value of $k$ indicates that a higher $ESE$ score leads to a more significant positive impact on the probability of success. This aligns well with the idea that a borrower’s positive $ESE$ practices contribute to a project success. In other words, higher $ESE$ scores indicate a better alignment with responsible practices. The constant $b$ represents a baseline/constant contribution of factors beyond an individual’s control, such as climatic and geographical conditions. Even if a borrower has a lower $ESE$ score, this constant term acknowledges that external factors can still influence the project’s success. It is essential to account for such factors in order to avoid unfairly penalizing borrowers for circumstances beyond their own control.
@@
The selection of the linear equation $e=k\cdot ESE+b$ is motivated by both the characteristics of the variables involved and its practical applicability. Notably, the non-negativity of $k$, $b$, and $ESE$ ensures that the resulting probability, denoted as $e$, maintains its intrinsic non-negative nature. As $ESE$ encompasses a continuous range from 0 to 100, it inherently aligns with this non-negative property. It is important to stress that the intuitive requirement for $e$ to lie between 0 and 1, which is a fundamental property of a probability, is elegantly met by setting $b=0$ and $k=\frac{1}{100}$. By adhering to these values, the probability of $e$ adheres to the established probability range. In addition, the model accommodates real-world scenarios by allowing for a non-zero baseline influence ($b>0$), often linked to exogenous factors. To this end, a modest value such as $b=0.3$ can be accommodated without violating the probability constraints. Notably, this adjustment merely necessitates setting $k=\frac{0.7}{100}$, thereby guaranteeing that $e$ remains within the bounds of a probability as it should be. While, at the same time, it also acknowledges that $e$ cannot span the interval from 0 to 0.3, this limitation is justified by underscoring $e$ as a representation of a probability of a high yield, which is inherently a positive entity in practical settings. Hence the rationale underlying the proposed equation effectively marries mathematical coherence with the nuanced demands from a real–world application.
@@
ASSUMPTION 3. The amount of loan secured by each borrower is denoted as $L$, while the amount for this borrower to be repaid is denoted as $w$. The unit of a selling price of the product is denoted as $p$, and the total income of a high production is denoted as $p\overline{Y}$, while the total income of a low production is denoted as $p\underline{Y}$.
@@
ASSUMPTION 4. High-yield borrowers repay their loans and low-yield borrowers pay back all the income they earn. As long as this group repays the total loans, this group can continue to secure subsequent loans. In the event that all borrowers are unsuccessful in their projects, the group will not be eligible for any future loans. This may result in a negative credit history for all members in the group, which could lead to further restrictions placed on them, such as being unable to use public transportation or purchase high-cost items. Therefore, if all borrowers fail, they may have to use all of their earnings to avoid receiving negative credit reports, even though their total earnings will be less than the total repayments.
@@
ASSUMPTION 5. To increase the successful probability $e$, the borrower incurs additional costs, such as purchasing green pesticides, advanced tools, etc. A convex disutility function, denoted by $C(e)$, is included in the specification of the total utility function in the proposed model.
@@ 2.2 Joint Liability Contracts with a Group Size of Two
Under a joint liability agreement, each farmer’s ex-ante expected profit $\pi_{i}$ can be written as
@@
We emphasize that, in this model, both farmers have a chance of achieving a successful harvest $\overline{Y}$ with a probability of $e^{2}$, which gives them a profit of $p\overline{Y}-w$. The probability of one farmer defaulting is given by $e(1-e)$, which means that farmer $i$ would receive his/her own surplus from a successful harvest minus the other farmer’s deficit, $p\overline{Y}+p\underline{Y}-2w$. If both farmers fail at the same time, they will pay all of their incomes to the lenders, which means that their profit will be zero. The detailed profit distributions are summarized conveniently in Table 1.
@@
One farmer succeeds, another farmer succeeds too. $e^{2}$ $p\overline{Y}-w$
One farmer succeeds, another farmer fails. $e(1-e)$ $p\overline{Y}+p\underline{Y}-2w$
One farmer fails, another farmer succeeds. $(1-e)e$ $p\underline{Y}-p\underline{Y}=0$
One farmer fails, another farmer fails. $(1-e)^{2}$ $p\underline{Y}-p\underline{Y}=0$
@@
Before setting the total objective (utility) function for the whole group in the model, we first list a number of exogenous constraints which we introduce to the corresponding optimization program.
CONSTRAINT 1. If the group wants a subsequent loan from the lender, the group’s total repayment amount must be affordable even in the worst–case scenario when everyone else’s project has failed except for him/hers:
@@
CONSTRAINT 2. Strategic default in a joint liability contract occurs when one party (or more parties) intentionally chooses (or choose) to default on his/her obligation (or their obligations), even though he/she has (or they have) the ability to pay back the loan. This is often done in order to gain a strategic advantage, such as to force the other parties to pay a greater share of the debt or renegotiate the terms of the contract. To avoid a strategic default for each successful borrower, the payoff of strategically defaulting cannot be larger than the payoff of repaying and being refinanced.
Assuming that the other borrower also succeeds in his/her project, this borrower will only be responsible for his/her own payment. In such a scenario the earnings plus the discounted cost of the next loan would at least be equal to the earnings of strategically defaulting.
@@
CONSTRAINT 3. If the other borrower fails in his/her project, then this borrower needs to pay for himself/herself as well as part of the other borrower’s repayment amount. Now, the earnings plus the discounted cost of the next loan would also at least equal the earnings of strategically defaulting.
@@
CONSTRAINT 4. The lender must be able to sustain the lending game, or at least break even, over the periods. Thus the expected repayment amount of each borrower has to be at least as large as $L(1+\epsilon)$, where $\epsilon$ is a risk-free rate. It is important to note that given that all constraints have been established under the assumption that the group will successfully repay the total loans to ensure their eligibility for subsequent loan rounds, we have not factored in the scenario where all borrowers within the group experience failure on the left hand side of Constraint 4.
@@
Since $p\underline{Y}-w<0$, it is clear that Constraint 3 is tighter than Constraint 2. In the ensuing analysis, we will only consider Constraint 1, Constraint 3, and Constraint 4. Firstly, $\pi_{i}$ and $w$ have a strictly negative relationship as can be observed from Equation (2.1). So, to maximize $\pi_{i}$, $w$ has to be set at a minimum value. From Constraint 4, the minimum value of $w$ should be set at
@@
. Next we substitute $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into Constraint 1, and rewrite the equation as:
@@
and the first loan ceiling bound can be expressed in Equation (2.8) as
@@
After substituting $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into Constraint 3, we obtain
@@
This allows us to obtain another bound for the loan ceiling as in Equation (2.11)
@@
Next we substitute $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into the objective function (which is equation (2.1) and replace $e=k\cdot E+b$. The ex-ante expected profit function of an individual borrower can now be rewritten as
@@
To simplify the problem even further, we set $C(e)=\frac{1}{2}ce^{2}$ (where the cost parameter $c$ is used to scale the total cost) as the disutility function. The disutility function $C(e)=\frac{1}{2}ce^{2}$ is a commonly used function in economics to represent the cost or disutility associated with effort or work. See e.g. clark1996satisfaction and kahneman1997back. The function is quadratic, meaning that as an effort increases, the cost or disutility of that effort increases at an increasing rate. In the context of an individual agricultural farmer, it is reasonable to use a disutility function to acknowledge that while some effort may be necessary, there is a point at which additional effort becomes more and more unpleasant or burdensome. For instance, a farmer may be willing to put in an extra effort to cultivate crops or tend to a livestock, but as the required effort increases, the farmer may experience fatigue, stress, and reduced productivity, which can negatively impact their overall well–being. Furthermore, the quadratic form of the disutility function is also supported by empirical evidence on the relationship between effort and well–being. It is reasonable to assume that farmers will experience a phenomenon known as a diminishing marginal utility of effort, which implies that the additional unit of well-being gained by exerting more effort decreases as the level of effort increases.
2.3 Main Results
Now that we have listed the assumptions and set-up for the model, we proceed to state a number of key results expressed conveniently in the form of propositions. These propositions provide useful insights into the optimal individual–ESE scores, optimal group size, and loan ceiling constraints. PROPOSITION 1 (Optimal individual–ESE score). Finding an optimal ESE score for both the lenders and borrowers in an individual ESE–joint liability model is critical because it can help address the challenge of promoting sustainability in agricultural production while ensuring access to credit for small-scale farmers. By incorporating environmental and social factors into credit evaluation systems, it can incentivize farmers to adopt sustainable practices that can lead to positive environmental outcomes and long–term economic benefits. At the same time lenders can mitigate the risks associated with lending to farmers by incorporating sustainability factors into their credit evaluation process. To find the optimal individual–$ESE$ score (noted in formulas as $E$), we take a first-order partial derivative w.r.t $E$ of the ex-ante expected payoff function to obtain the following result
@@ 2.3 Main Results
After rearranging terms, we obtain an optimal individual-$ESE$ score (where $E$ is between 0 and 100) as
@@
PROPOSITION 2 (Collapsing two constraints on loan ceiling to a single constraint). The two loan ceiling constraints (Constraint 1 and Constraint 3) arise from analyzing the success or failure of the other co–loaners assuming their own success in the project, but most of the previous studies in this area have not explored the connection between these constraints. Notably, in rezaei2017optimal, they found that a feasible range of loan ceiling depends on the borrower’s discount rate $\delta$. Nevertheless, it is possible to reduce these two constraints to a single constraint in our model. Recall the two constraints on the loan ceiling in Equation (2.8) and Equation (2.11), labelled respectively as $L_{1}$ and $L_{2}$.
@@
Next we show that $L_{1}$ is always larger than $L_{2}$, which means that we can only use $L_{2}$ as a final loan ceiling bound. It simplifies the problem and potentially makes it easier to analyze and solve the model. In addition, $L_{2}$ takes $\delta$ into consideration, which may include more information and additional key factors. This conclusion is illustrated in Figure 1.
@@
Proof: If $L_{1}\leqslant L_{2}$, then
@@
After rearranging terms, we obtain
@@
From the above results, if $L_{1}\leqslant L_{2}$, the discount factor $\delta$ is always greater than 1, which leads to a contradiction. This leads us to conclude that $L_{1}>L_{2}$. PROPOSITION 3 (A higher individual–ESE score leads to a higher loan ceiling). Next we examine the relationship between the individual–ESE score and the loan ceiling. Intuitively a borrower with a high-quality ESE score should be eligible for a higher loan amount, and therefore the loan ceiling should also be higher. Since $L_{1}$ is always greater than $L_{2}$, we will use $L_{2}$ to determine the loan ceiling, which defines the maximum value of $L_{2}$ to be the loan ceiling.
@@
Then we take a partial derivative of $L_{2}$ w.r.t. $E$ to obtain
@@
It is clear that the denominator is greater than 0. The numerator is also greater than 0, which can be illustrated by writing out the following result
@@
A borrower with a higher optimal ESE score, which means he/she has a better overall performance, will be perceived as less risky by the lender. As a result the borrower will be offered a higher loan ceiling, which is the maximum amount of loan that a lender is willing to provide to a borrower. In turn, if a borrower wants to obtain a larger loan amount, he/she may be motivated to improve his/her ESE performance.